in the binomial expantion (a+bx)^(-3)=1/8+9/8x+...... then the value of a & b is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear manan

(a+bx)^(-3)=1/8+9/8x+......

u can solve it by using binomial theorem

 

here just put x= 0

(a)^(-3)=1/8

 a =2

now differentiate

-3(a+bx)-4(b) = 9/8 + ........

 

now put x=0

-3b/a4 =9/8

-3b/16 =9/8

 b =-6

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin